What is the degree of unsaturation for the following molecule? HO O OOOOO
Answer #1

Degree of unsaturation = 2c +2+N-H-X - 2 c= Carbon X= Nor of halogen H=hydrogen IN-H=1-8) -71 ► 2x8+2+7-1 2 HO = 10/2=5